英語 (English)

主題別英語 (Thematic English) (薬 (薬 A1 年))

早内-プリングルジュディス · 非常勤講師 1 単位後期火 5 ・ 6

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

【授業の目的】To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary,grammatical ability, reading ability and fluency. Various media forms will be used in this class.

【授業の概要】 Speaking, reading, presentation Notes: 1. Students are expected to speak and write only in English during class. They are expected to work not only in class, but outside of class as well. Homework is required, and they are expected to complete it even if they are absent from class. 2. The following class schedule is tentative and only shows the general contents which will be adjusted due to student input and other factors. There may be other modifications based on class size, etc.

【キーワード】 Thematic English， fluency， Twilight 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】In today’s globalized society, English has become the common

language of communication and is necessary for gaining more knowledge of world events, and to process information. This class requires students to put their skills into practice, and while doing so, increase fluency and confidence in using English while also increasing vocabulary and polishing grammatical skills.

【授業の計画】 1. General Class Guidance Introduction; Survey

2. Twilight: Introduction to the movie, book & cultural background

3. Movie Scene 1; listening & reading activities

4. Movie Scene 2; listening & writing activities

5. Movie Scene 3; listening & role play

6. Movie Scene 4; listening & discussion

7. Movie Scene 5; listening & group activity

8. Mid-term evaluation; presentation

9. Movie Scene 6; listening & writing activities 10. Movie Scene 7; listening & role play

11. Movie Scene 8; listening & discussion

12. Movie Scene 9; listening & interview

13. Movie Scene 10; listening & reading

14. Movie Scene 11; listening & discussion

15. Final Evaluation; Presentations Week 1

16. Final Presentations Week 2

【教科書】 Class handouts and other materials will be provided as necessary. Some

content will also be student-generated. 【参考書等】 Twilight by Stephenie Meyer; Twilight: The Movie 【成績評価の方法】 Grading will be as follows: Class Attendance and participation (including class activities, homework, etc):50% Mid-term and final evaluations: 20% Skits/Role plays/mini-presentations: 30%

【再試験の有無】 Allowable if student has attendee more than 2/3 of the classes and

has a score of 50% or higher. 【受講者へのメッセージ】 [注意]

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221173 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 早内-プリングル (juditheph@yahoo.com) ���� (オフィスアワー: By appointment)

English

Thematic English (薬 (薬 A1 年))

Judith Hayauchi · Part-time Lecturer 1 unit 後期火 5 ・ 6

(平成 19 年度以前の授業科目: 『主題別英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (2)』 )

Target⟩ To give students opportunities to develop their English abilities through the exploration and discussion of a variety of themes. Students will develop their real-life communicative skills while expanding their vocabulary,grammatical ability, reading ability and fluency. Various media forms will be used in this class.

Outline⟩ Speaking, reading, presentation Notes: 1. Students are expected to speak and write only in English during class. They are expected to work not only in class, but outside of class as well. Homework is required, and they are expected to complete it even if they are absent from class. 2. The following class schedule is tentative and only shows the general contents which will be adjusted due to student input and other factors. There may be other modifications based on class size, etc.

�������⟩ Thematic English, fluency, Twilight Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩ In today’s globalized society, English has become the common language of

communication and is necessary for gaining more knowledge of world events, and to process information. This class requires students to put their skills into practice, and while doing so, increase fluency and confidence in using English while also increasing vocabulary and polishing grammatical skills.

Schedule⟩

1. General Class Guidance Introduction; Survey

2. Twilight: Introduction to the movie, book & cultural background

3. Movie Scene 1; listening & reading activities

4. Movie Scene 2; listening & writing activities

5. Movie Scene 3; listening & role play

6. Movie Scene 4; listening & discussion

7. Movie Scene 5; listening & group activity

8. Mid-term evaluation; presentation

9. Movie Scene 6; listening & writing activities 10. Movie Scene 7; listening & role play

11. Movie Scene 8; listening & discussion

12. Movie Scene 9; listening & interview

13. Movie Scene 10; listening & reading

14. Movie Scene 11; listening & discussion

15. Final Evaluation; Presentations Week 1

16. Final Presentations Week 2

Textbook⟩ Class handouts and other materials will be provided as necessary. Some

content will also be student-generated. Reference⟩ Twilight by Stephenie Meyer; Twilight: The Movie Evaluation Criteria⟩ Grading will be as follows: Class Attendance and participation (including class activities, homework, etc):50% Mid-term and final evaluations: 20% Skits/Role plays/mini-presentations: 30%

Re-evaluation⟩ Allowable if student has attendee more than 2/3 of the classes

and has a score of 50% or higher. Message⟩ [注意]

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=221173 Contact (Office-Hour, Room, E-mail)⟩

⇒ Hayauchi (juditheph@yahoo.com) ���� (Office Hour: By appointment)